Verify a – (-b) = a + b for the following values of a and 6.
(i) a = 21, b = 18
(ii) a = 118, b = 125
(iii) a = 75, b = 84
(iv) a= 28, 6 = 11

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) a – (-b) = a + b LHS = 21 – (-18) = 21 + 18 = 39 RHS = 21 + 18 = 39 LHS = RHS Hence, verified. (ii) a – (-b) = a + b LHS = 118 – (-125) = 118 + 125 = 243 RHS = 118 + 125 = 243 LHS = RHS Hence, verified. (iii) a – (-b) = a + b LHS = 75 – (-84) = 75 + 84 = 159 RHS = 75 + 84 = 159 LHS = RHS Hence, verified. (iv) a – (-b) = a + b LHS = 28 – (-11) = 28 + 11 = 39 RHS = 28 + 11 = 28 + 11 = 39 LHS = RHS Hence, verified